mikes age is three -fourths of marys age. in five years mikes age will be 7/9 marys age . how old is mike now?
Answer by venugopalramana(3286)

mikes age ( M SAY)is three -fourths of marys(R SAY) age.
M =(3/4)*R NOW

in five years mikes age will be 7/9 marys age . how old is mike now?
IN 5 YEARS MIKE WILL BE = M+5 YRS
IN 5 YRS MARY WILL BE =R+5 YRS
(M+5)=(7/9)(R+5)
9(M+5)=7(R+5)
9M+45=7R+35
7R-9M=45-35=10
7R-9*(3/4)*R=10
R[7-27/4]=10
R(28-27)/4=10
R=40 YRS.
M=(3/4)*40=30 YRS.